### v1

@@ Abstract
We introduce multi-jet merging for deep inelastic scattering in the Vincia parton shower in the Monte Carlo event generator Pythia 8. Merging combines event samples of different parton multiplicities with logarithmically enhanced parton-shower radiation. We consider up to five outgoing partons using two different merging algorithms. We vary the relevant scale choices and compare to experimental data by H1 collaboration at the HERA collider. Results show that multi-jet merging improves description to data, especially for low virtuality events, and that jet cross-sections converge when the partonic multiplicity exceeds the number of reconstructed jets.
@@ 1 Introduction
Matrix-element calculations, however, get increasingly complicated and computationally demanding due to the intrinsic factorial scaling of the number of diagrams with the number of particles. Under the assumption that the radiation of additional particles appears at small relative angles (collinear) or at small relative energies (soft), radiative corrections can be modelled by parton showers, which evolve partons to lower energies through successive branchings bierlich2022comprehensive . In traditional parton showers, these branchings are sampled according to the DGLAP-evolution equations Dokshitzer:1977sg ; Gribov:1972ri ; Lipatov:1974qm ; Altarelli:1977zs , which resum large logarithms arising from collinear emissions. Modern parton showers typically use some form of colour dipoles Sjostrand:2004ef ; Nagy:2005aa ; Schumann:2007mg ; Winter:2007ye ; Giele:2007di ; Nagy:2007ty ; Hoche:2015sya ; Brooks_2020 , with angular-ordered showers providing a notable exception Gieseke:2003rz ; Platzer:2009jq .
@@
In this study, we implement a multi-jet merging algorithm for deep inelastic scattering (DIS) events in Pythia bierlich2022comprehensive , utilising the Vincia antenna shower Brooks_2020 . We employ the two merging algorithms mentioned above, CKKW-L and UMEPS, and consider two different options for the merging scale combined with appropriate factorization- and renormalization-scale choices. Multi-jet merging in DIS has previously been studied in the context of the Sherpa event generator in Carli_2010 and formed the basis of a NNLO matched calculation in H_che_2018 . In earlier works ZEUS:1998uhq ; ZEUS:1999nrh ; Carli_2010 it has been shown that especially at low virtualities, a simplistic MC treatment of parton showers does not provide a good description of (multi-)jet final states. In light of this, the presented merging implementation will ensure accurate predictions of DIS jet measurements for the forthcoming Electron-Ion Collider (EIC) ABDULKHALEK2022122447 and can be used to revisit diffractive dijet production cross sections measured at HERA to further study the observed breakdown of factorization Helenius:2019gbd .
@@ 2.1 Parton showers
Parton shower algorithms are a crucial part of Monte Carlo event generation. They make use of universal splitting functions to describe the evolution of a scattering event in some energy scale and are therefore intrinsically process-independent. The evolution is governed by the parton shower evolution variable, which, in broad terms, represents the hardness of the event. In doing so, logarithmically enhanced terms are resummed. The formal accuracy of the PS resummation may be expressed in terms of a power series in logarithms, as leading-logarithmic (LL), next-to-leading-logarithmic (NLL) and so on. In the present study we can generally only claim formal LL correctness, although there are recent developments on NLL-consistent parton shower algorithms Bewick:2019rbu ; Dasgupta:2020fwr ; Forshaw:2020wrq ; Herren:2022jej ; Bewick:2021nhc ; vanBeekveld:2022ukn ; vanBeekveld:2022zhl ; Assi:2023rbu ; Hoche:2024dee ; Preuss:2024vyu and beyond FerrarioRavasio:2023kyg ; vanBeekveld:2024wws , also in the context of DIS processes vanBeekveld:2023chs .
The default parton shower algorithm in Pythia is the transverse-momentum-ordered “simple shower” algorithm Sjostrand:2004ef . In the simple shower, branching probabilities are obtained from DGLAP splitting kernels $\hat{P}_{a\rightarrow bc}(z)$. The probability for a parton $a$ to split into partons $b$ and $c$ can thus be formally written as
@@
where $t$ is the parton-shower evolution variable and $z$ the energy fraction of parton $b$. Parton showers employ the veto algorithm to produce emissions. An extension of Pythia’s simple shower to introduce dipole-antenna-like initial-final radiation patterns into the evolution has been given in Cabouat:2017rzi , which includes a recoil handling appropriate for coherent branchings in DIS.
@@
This ensures that, at least at leading colour, Vincia faithfully accounts for the coherent suppression of soft-gluon radiation in initial-final (IF) dipoles, as encountered in DIS and VBF processes Cabouat:2017rzi ; Hoche:2021mkv . Parton branchings are simulated in Vincia as an interleaved sequence ordered in transverse momentum defined as
@@ 2.2 Multi-jet merging
the no-emission probability of state $\mathcal{S}_{n}$ between scales $t_{1}$ and $t_{2}$. The following sections are written with simplified notation, considering only final state radiation (FSR) of parton showers while omitting initial-state radiation (ISR), and taking implicit integrals in $z$. If we attach the parton shower to the core process $B_{0}$, the exclusive parton shower cross sections become
$\frac{\mathrm{d}\sigma_{0}^{\mathrm{PS}}}{\mathrm{d}\Phi_{0}}$ $=\mathrm{B}_{0}\Pi_{\mathcal{S}_{0}}(t_{0},t_{c})$
$\frac{\mathrm{d}\sigma_{1}^{\mathrm{PS}}}{\mathrm{d}\Phi_{0}}$ $=\int_{t_{c}}^{t_{0}}\mathrm{d}t\mathrm{B}_{0}\Pi_{\mathcal{S}_{0}}(t_{0},t)\mathcal{P}_{\mathcal{S}_{0}}(t)\Pi_{\mathcal{S}_{1}}(t,t_{c})$
@@
where $t_{c}\sim\Lambda_{\mathrm{QCD}}^{2}$ is the parton shower cutoff.
With this notation at hand, let us write the effect of a general merging algorithm to the exclusive $+2$-jet cross section. Assume that a matrix element generator generates events according to the cross sections $B_{n}$ with all partons above the merging scale, $t>t_{\mathrm{MS}}$, and we attach the parton shower to these samples with emissions below $t_{\mathrm{MS}}$. The integrated exclusive $+2$-jet cross section, that is, the cross section for exactly two additional jets, can then be written as
@@
$w^{\alpha_{\mathrm{S}}}$ $=\prod_{i=1}^{n}\frac{\alpha_{\mathrm{S}}(t_{i})}{\alpha_{\mathrm{S}}(\mu_{\mathrm{R}})}$ $w^{\mathrm{Sudakov}}$ $=\Pi_{\mathcal{S}_{n}}(t_{n},t_{\mathrm{MS}})\,\prod_{i=1}^{n}\Pi_{\mathcal{S}_{i-1}}(t_{i-1},t_{i}).$
@@
    Perform trial emissions from states $0\leq i<n$. Reject the event if the generated emission is above $t_{\mathrm{MS}}$. This step has the same effect as multiplying with the corresponding no-branching probability. Reweight the event with PDF ratios and $\alpha_{\mathrm{S}}$ ratios.
@@
    Accept the event and continue the parton-shower cascade from $t_{n}$ to $t_{c}$ if
    • $n<N$, the jet multiplicity is not the highest and the first emission is below $t_{\mathrm{MS}}$ or
    • $n=N$, the jet multiplicity is the highest.
    Otherwise reject the whole event.
@@
$\frac{\mathrm{d}\sigma_{0}^{\mathrm{ex}}}{\mathrm{d}\Phi_{0}}$ $=\mathrm{B}_{0}\Pi_{\mathcal{S}_{0}}(t_{0},t_{\mathrm{MS}})=w_{0}\mathrm{B}_{0}$
$\frac{\mathrm{d}\sigma_{1}^{\mathrm{ex}}}{\mathrm{d}\Phi_{0}}$ $=\frac{f_{1}(x_{1},t_{1})}{f_{0}(x_{0},t_{1})}\frac{\alpha_{\mathrm{S}}(t_{1})}{\alpha_{\mathrm{S}}(\mu_{\mathrm{R}})}\Pi_{\mathcal{S}_{0}}(t_{0},t_{1})\Pi_{\mathcal{S}_{1}}(t_{1},t_{\mathrm{MS}})\,\mathrm{B}_{1}=w_{1}\mathrm{B}_{1}$
$\frac{\mathrm{d}\sigma_{2}^{\mathrm{in}}}{\mathrm{d}\Phi_{0}}$ $=\frac{f_{1}(x_{1},t_{1})}{f_{0}(x_{0},t_{1})}\frac{\alpha_{\mathrm{S}}(t_{1})}{\alpha_{\mathrm{S}}(\mu_{\mathrm{R}})}\frac{f_{2}(x_{2},t_{2})}{f_{1}(x_{1},t_{2})}\frac{\alpha_{\mathrm{S}}(t_{2})}{\alpha_{\mathrm{S}}(\mu_{\mathrm{R}})}\Pi_{\mathcal{S}_{0}}(t_{0},t_{1})\Pi_{\mathcal{S}_{1}}(t_{1},t_{2})\Pi_{\mathcal{S}_{2}}(t_{2},t_{\mathrm{MS}})\,\mathrm{B}_{2}=w_{2}\mathrm{B}_{2}.$
The sum of these cross sections is the $+2$-jet inclusive CKKW-L merged cross section,
$\sigma_{2}^{\mathrm{CKKW-L}}=\int\mathrm{d}\Phi_{0}\int\mathrm{d}t\Bigg{(}w_{0}\mathrm{B}_{0}+\int\mathrm{d}t_{1}w_{1}\mathrm{B}_{1}+\int\int\mathrm{d}t_{1}\mathrm{d}t_{2}w_{2}\mathrm{B}_{2}\Bigg{)}$ (12)
which can reproduce the Born-level cross section only if certain conditions are met, one being that the parton-shower expansion exactly reproduces the matrix element contributions for the first two emissions. Further conditions arise for multiplicities with $n>1$, for example identical phase space coverage no-branching probabilities. The task of generating PS emissions according to the ME is possible to be done process-by-process, and has potential for a future study on Pythia DIS improvements. These so called matrix-element corrections have been implemented for various parton shower algorithms and processes Giele:2007di ; Bengtsson:1986hr ; Fischer:2017yja .
@@
$\frac{\mathrm{d}\sigma_{n}^{\mathrm{in}}}{\mathrm{d}\Phi_{0}}$ $=w^{\prime}_{n}\mathrm{B}_{n}=\widehat{\mathrm{B}}_{n},$ $\frac{\mathrm{d}\sigma_{n\rightarrow m}^{\mathrm{in}}}{\mathrm{d}\Phi_{0}}=\int\mathrm{d}^{n-m}\Phi\,w^{\prime}_{n}\mathrm{B}_{n}=\int\widehat{\mathrm{B}}_{n\rightarrow m},$
so that the UMEPS cross section can be written as a sum of positive-weight $n$-jet samples and negative-weight $(n-m)$-jet samples. In the case of $+2$-jet merging, the inclusive cross section becomes
@@ 2.4 Merging in DIS
A fixed merging scale may suffice for many processes where the highest energy scale is known, such as Drell-Yan or electron-positron-annihilation, where it is the center-of-mass energy of the incoming interacting particles Carli_2010 ; Lavesson:2005xu ; Lonnblad:2011xx . We denote $\tilde{t}_{\mathrm{MS}}$ as the fixed-value merging scale which separates low-transverse-momenta jets from hard jets. It serves as the cut in the merging algorithm, in principle assigning vetoes to matrix-element events below this scale. As discussed above, the hardest scale in a DIS event is not unique, and this handling with the usual choice of $\mu_{\mathrm{F}}^{2}=Q^{2}$ might miss a large portion of the low-$Q^{2}$, high-$p_{\mathrm{T}}$ phase space. A re-definition of the merging scale aims to provide a similar effect as switching to a more suitable factorization scale. To account for the contributions of these processes, a dynamic merging scale is defined as in the study Carli_2010 , as shown by the following equation:

### v3

@@ Abstract
We introduce multi-jet merging for deep inelastic scattering in the Vincia parton shower in the Monte Carlo event generator Pythia 8. Merging combines event samples of different parton multiplicities with logarithmically enhanced parton-shower radiation. We consider up to five outgoing partons using two different merging algorithms. We vary the relevant scale choices and compare to experimental data by the H1 collaboration at the HERA collider. Results show that multi-jet merging improves the description to data, especially for low virtuality events, and that jet cross-sections converge when the partonic multiplicity exceeds the number of reconstructed jets.
Keywords:
Deep Inelastic Scattering or Small-x Physics, Jets and Jet Substructure, Parton Shower
arxiv: 2410.20950
@@ 1 Introduction
Matrix-element calculations, however, get increasingly complicated and computationally demanding due to the intrinsic factorial scaling of the number of diagrams with the number of particles. Under the assumption that the radiation of additional particles appears at small relative angles (collinear) or at small relative energies (soft), radiative corrections can be modelled by parton showers, which evolve partons to lower energies through successive branchings. In traditional parton showers, these branchings are sampled according to the DGLAP-evolution equations Dokshitzer:1977sg ; Gribov:1972ri ; Lipatov:1974qm ; Altarelli:1977zs , which resum large logarithms arising from collinear emissions. Modern parton showers typically use some form of colour dipoles Sjostrand:2004ef ; Nagy:2005aa ; Schumann:2007mg ; Winter:2007ye ; Giele:2007di ; Nagy:2007ty ; Hoche:2015sya ; Brooks_2020 ; Platzer:2009jq ; Dasgupta:2020fwr ; Forshaw:2020wrq ; Preuss:2024vyu , with angular-ordered showers providing a notable exception Gieseke:2003rz .
@@
In this study, we implement a multi-jet merging algorithm for deep inelastic scattering (DIS) events in Pythia bierlich2022comprehensive , utilising the Vincia antenna shower Brooks_2020 . We employ the two merging algorithms mentioned above, CKKW-L and UMEPS, and consider two different options for the merging scale combined with appropriate factorization- and renormalization-scale choices. Multi-jet merging in DIS has previously been studied in the context of the Sherpa event generator in Carli_2010 which formed the basis of a NNLO matched calculation in H_che_2018 and of NLO multi-jet merging in knobbe2023nnlonll . A number of dedicated Powheg-Box NLO+PS related works have recently been presented in Banfi_2024 ; Borsa:2024rmh ; Buonocore:2024pdv ; FerrarioRavasio:2024kem . In earlier works ZEUS:1998uhq ; ZEUS:1999nrh ; Carli_2010 it has been shown that especially at low virtualities, a simplistic MC treatment of parton showers does not provide a good description of (multi-)jet final states. In light of this, the presented merging implementation will ensure accurate predictions of DIS jet measurements for the forthcoming Electron-Ion Collider (EIC) ABDULKHALEK2022122447 and can be used to revisit diffractive dijet production cross sections measured at HERA to further study the observed breakdown of factorization Helenius:2019gbd .
@@ 2.1 Parton showers
Parton shower algorithms are a crucial part of Monte Carlo event generation. They make use of universal splitting functions to describe the evolution of a scattering event in some energy scale and are therefore intrinsically process-independent. The evolution is governed by the parton-shower evolution variable, which, in broad terms, represents the hardness of the event. In doing so, logarithmically enhanced terms are resummed. The formal accuracy of the PS resummation may be expressed in terms of a power series in logarithms, as leading-logarithmic (LL), next-to-leading-logarithmic (NLL) and so on. In the present study we can generally only claim formal LL correctness, although there are recent developments on NLL-consistent parton shower algorithms Bewick:2019rbu ; Dasgupta:2020fwr ; Forshaw:2020wrq ; Herren:2022jej ; Bewick:2021nhc ; vanBeekveld:2022ukn ; vanBeekveld:2022zhl ; Assi:2023rbu ; Hoche:2024dee ; Preuss:2024vyu and beyond FerrarioRavasio:2023kyg ; vanBeekveld:2024wws , also in the context of DIS processes vanBeekveld:2023chs .
The default parton shower algorithm in Pythia is the transverse-momentum-ordered “simple shower” algorithm Sjostrand:2004ef . Given that the simple shower is known to not faithfully reproduce coherent branchings in initial-final dipoles Jager:2014vna ; Rauch:2016upa ; Jager:2020hkz ; Covarelli:2021gyz ; Hoche:2021mkv , it is in particular not suitable for DIS topologies. An extension to introduce dipole-antenna-like initial-final radiation patterns into the evolution has been given in Cabouat:2017rzi , which includes a recoil handling appropriate for coherent branchings in DIS. In the simple shower, branching probabilities are obtained from DGLAP splitting kernels $\hat{P}_{a\rightarrow bc}(z)$. The probability for a parton $a$ to split into partons $b$ and $c$ can thus be formally written as
@@
where $t$ is the parton-shower evolution variable and $z$ the energy fraction of parton $b$. Parton showers employ the veto algorithm to produce emissions.
@@
This ensures that, at least at leading colour, Vincia faithfully accounts for the coherent suppression of soft-gluon radiation in initial-final (IF) dipoles, as encountered in DIS and vector-boson-fusion (VBF) processes Cabouat:2017rzi ; Hoche:2021mkv . This has also been addressed in other studies regarding to VBF Jager:2020hkz ; Ballestrero:2018anz . Parton branchings are simulated in Vincia as an interleaved sequence ordered in transverse momentum defined as
@@ 2.2 Multi-jet merging
the no-emission probability of state $\mathcal{S}_{n}$ between scales $t_{1}$ and $t_{2}$. The following sections are written with simplified notation, considering only final-state radiation (FSR) of parton showers while omitting initial-state radiation (ISR), and taking implicit integrals in $z$. If we attach the parton shower to the core process $\mathrm{B}_{0}$, the exclusive parton-shower cross-sections become
$\frac{\mathrm{d}\sigma_{0}^{\mathrm{PS}}}{\mathrm{d}\Phi_{0}}$ $=\mathrm{B}_{0}\Pi_{\mathcal{S}_{0}}(t_{0},t_{c})$ (9)
$\frac{\mathrm{d}\sigma_{1}^{\mathrm{PS}}}{\mathrm{d}\Phi_{0}}$ $=\int_{t_{c}}^{t_{0}}\mathrm{d}t\mathrm{B}_{0}\Pi_{\mathcal{S}_{0}}(t_{0},t)\mathcal{P}_{\mathcal{S}_{0}}(t)\Pi_{\mathcal{S}_{1}}(t,t_{c})$ (10)
@@
where $t_{c}\sim\Lambda_{\mathrm{QCD}}^{2}$ is the parton-shower cutoff.
With this notation at hand, let us write the effect of a general merging algorithm to the exclusive $+2$-jet cross section. Assume that a matrix element generator generates events according to the cross sections $\mathrm{B}_{n}$ with all partons above the merging scale, $t>t_{\mathrm{MS}}$, and we attach the parton shower to these samples with emissions below $t_{\mathrm{MS}}$. The integrated exclusive $+2$-jet cross section, that is, the cross section for exactly two additional jets, can then be written as
@@
$w^{\alpha_{\mathrm{S}}}$ $=\prod_{i=1}^{n}\frac{\alpha_{\mathrm{S}}(t_{i})}{\alpha_{\mathrm{S}}(\mu_{\mathrm{R}})}$ $w^{\mathrm{Sudakov}}$ $=\Pi_{\mathcal{S}_{n}}(t_{n},t_{\mathrm{MS}})\,\prod_{i=1}^{n}\Pi_{\mathcal{S}_{i-1}}(t_{i-1},t_{i}).$ (14)
@@
    Perform trial emissions from states $0\leq i<n$. Reject the event if the generated emission is above $t_{i+1}$. This step has the same effect as multiplying with the corresponding no-branching probabilities. Reweight the event with PDF ratios and $\alpha_{\mathrm{S}}$ ratios.
@@
    Start the parton-shower cascade from $t_{n}$.
    • if $n<N$: reject the event if the first emission is above $t_{\mathrm{MS}}$, otherwise accept it;
    • if $n=N$: accept the event.
    If the event has been accepted, run the shower down to $t_{\mathrm{c}}$.
@@
$\frac{\mathrm{d}\sigma_{0}^{\mathrm{ex}}}{\mathrm{d}\Phi_{0}}$ $=\mathrm{B}_{0}\Pi_{\mathcal{S}_{0}}(t_{0},t_{\mathrm{MS}})=w_{0}\mathrm{B}_{0}$ (15)
$\frac{\mathrm{d}\sigma_{1}^{\mathrm{ex}}}{\mathrm{d}\Phi_{0}}$ $=\frac{f_{1}(x_{1},t_{1})}{f_{0}(x_{0},t_{1})}\frac{\alpha_{\mathrm{S}}(t_{1})}{\alpha_{\mathrm{S}}(\mu_{\mathrm{R}})}\Pi_{\mathcal{S}_{0}}(t_{0},t_{1})\Pi_{\mathcal{S}_{1}}(t_{1},t_{\mathrm{MS}})\,\mathrm{B}_{1}=w_{1}\mathrm{B}_{1}$ (16)
$\frac{\mathrm{d}\sigma_{2}^{\mathrm{in}}}{\mathrm{d}\Phi_{0}}$ $=\frac{f_{1}(x_{1},t_{1})}{f_{0}(x_{0},t_{1})}\frac{\alpha_{\mathrm{S}}(t_{1})}{\alpha_{\mathrm{S}}(\mu_{\mathrm{R}})}\frac{f_{2}(x_{2},t_{2})}{f_{1}(x_{1},t_{2})}\frac{\alpha_{\mathrm{S}}(t_{2})}{\alpha_{\mathrm{S}}(\mu_{\mathrm{R}})}\Pi_{\mathcal{S}_{0}}(t_{0},t_{1})\Pi_{\mathcal{S}_{1}}(t_{1},t_{2})\,\mathrm{B}_{2}\equiv w^{\prime}_{2}\mathrm{B}_{2},$ (17)
where $w_{n}$ differs from $w^{\prime}_{n}$ only by a factor of $\Pi_{\mathcal{S}_{n}}(t_{n},t_{\mathrm{MS}})$. The sum of these cross sections is the $+2$-jet inclusive CKKW-L merged cross section,
$\sigma_{2}^{\mathrm{CKKW-L}}=\int\mathrm{d}\Phi_{0}\int\mathrm{d}t\Bigg{(}w_{0}\mathrm{B}_{0}+\int\mathrm{d}t_{1}w_{1}\mathrm{B}_{1}+\int\int\mathrm{d}t_{1}\mathrm{d}t_{2}w^{\prime}_{2}\mathrm{B}_{2}\Bigg{)}$ (18)
which can reproduce the Born-level cross section only if certain conditions are met, one being that the parton-shower expansion exactly reproduces the matrix element contributions for the first two emissions. Further conditions arise for multiplicities with $n>1$, for example identical phase space coverage and no-branching probabilities. The task of generating PS emissions according to the ME is possible to be done process-by-process, and has potential for a future study on Pythia DIS improvements. These so called matrix-element corrections have been implemented for various parton shower algorithms and processes Giele:2007di ; Bengtsson:1986hr ; Fischer:2017yja .
@@
$\frac{\mathrm{d}\sigma_{n}^{\mathrm{in}}}{\mathrm{d}\Phi_{0}}$ $=w^{\prime}_{n}\mathrm{B}_{n}\equiv\widehat{\mathrm{B}}_{n},$ $\frac{\mathrm{d}\sigma_{n\rightarrow m}^{\mathrm{in}}}{\mathrm{d}\Phi_{0}}=\int\mathrm{d}^{n-m}\Phi\,w^{\prime}_{n}\mathrm{B}_{n}\equiv\int\widehat{\mathrm{B}}_{n\rightarrow m},$ (19)
so that the UMEPS cross section can be written as a sum of positive-weight $n$-jet samples and negative-weight $(n-m)$-jet samples. Here the expression $\int\mathrm{d}^{n-m}\Phi\,w^{\prime}_{n}\mathrm{B}_{n}$ contains integrations of states $\mathcal{S}_{n-1},\mathcal{S}_{n-2},...,\mathcal{S}_{m+2},\mathcal{S}_{m+1}$. In this way it is possible to induce resummation in $m$-jet cross sections by integrating over the ($n>m$)-jet states $n-m$ times. In practice, this integration is the replacement of state $\mathcal{S}_{n}$ with state $\mathcal{S}_{n-1}$ while keeping the weight of the original state. These states are already obtained while constructing PS histories in the CKKW-L procedure. In the case of $+2$-jet merging, the inclusive cross section becomes
@@ 2.4 Merging in DIS
A fixed merging scale is sufficient for processes that can be characterised by a single scale, such as Drell-Yan or electron-positron-annihilation where the virtuality of the intermediate boson serves as a characteristic scale. This approach is insufficient whenever multiple scales are present, such as in DIS where the transverse momentum in the Breit frame and the virtuality of the exchanged boson characterize intrinsically different phase-space regions. Notably, VBF-processes, sometimes viewed as “double-DIS” processes Han:1992hr , are sufficiently well described by scale choices considering the transverse masses of the Boson and Born-level jets Buckley:2021gfw ; Hoche:2021mkv .
We denote $\tilde{t}_{\mathrm{MS}}$ as the fixed-value merging scale which separates low-transverse-momenta jets from hard jets. It serves as the cut in the merging algorithm, in principle assigning vetoes to matrix-element events below this scale. As discussed above, the hardest scale in a DIS event is not unique, and this handling with the usual choice of $\mu_{\mathrm{F}}^{2}=Q^{2}$ might miss a large portion of the low-$Q^{2}$, high-$p_{\mathrm{T}}$ phase space. A re-definition of the merging scale aims to provide a similar effect as switching to a more suitable factorization scale. To account for the contributions of these processes, a dynamic merging scale is defined as in the study Carli_2010 , as shown by the following equation: